OpenStudy (anonymous):

The area of a circular trampoline is 112.07 square feet. What is the radius of the trampoline? Round to the nearest hundredth. Answer 3.37 feet 10.59 feet 5.97 feet 35.67 feet

OpenStudy (anonymous):

use A=pi*radius^2 and solve for the radius: A=pi*r^2 112.07 = pi*r^2 divide both sides by pi: 112.07/pi = r^2 take the square root both sides: sqrt(112.07/pi) = r

OpenStudy (anonymous):

i got 3.24

OpenStudy (anonymous):

33.24

OpenStudy (anonymous):

actually i got 18.75